Could u please tell me how to find ans to this question


Q)the coordinates of foot of perpendicular drawn from a point A(1,0,3) to join of the points B(4,7,1) and c(3,5,3) are?

Grade:11

1 Answers

apoorva sagarwal
11 years ago

equation of li

 

 

 

AD is  l  to BC 

===>      (k+3)1+(2k+7)2+(-2k-2)-2 =0

===>      k+3+4k+14+4k+4=0

===>      9k+21=0

===>      k= -7/3

D  is  (-7/3+4,-14/3+7,14/3-2)

this is the foot of perpendicular